OpenStudy (anonymous):

Solve the system by elimination. x + y = 3 x – y = 5

OpenStudy (texaschic101):

what happens when you add these equations together ?

OpenStudy (texaschic101):

try adding them and what do you get ?

OpenStudy (texaschic101):

if you add them together, the y's will cancel out...then you can solve for x

OpenStudy (anonymous):

X+y=5 X-y=-3 Added together I get: 2x = 2 <--- (+y + -y = 0, so they cancel away... 5 + -3 = 2) Divide both sides by two to get X = 1 Use this answer in either equation to find y. I will use the first: X+y=5 (1)+y=5 Subtract 1 on both sides to get Y=4 Your answer should be a point (x, y)... So (1, 4

OpenStudy (texaschic101):

is that a -3 or just 3 in the first equation ??

OpenStudy (texaschic101):

oops...2nd equation

OpenStudy (texaschic101):

you got different problems ? Which one is the problem you need help with

OpenStudy (texaschic101):

you did the last problem correctly :)

OpenStudy (anonymous):

Coplanar lines are parallel if and only if they do not intersect. Write a definition based on the given biconditional. A. Parallel lines are coplanar. B. Parallel lines do not intersect. C. If coplanar lines are parallel, then they do not intersect. D. Parallel lines are coplanar lines that do not intersect.

OpenStudy (anonymous):

can someone help ?

OpenStudy (anonymous):

pleaseeeeeeeee

OpenStudy (texaschic101):

I would say C